An electrician works thrice as fast as his apprentice. Together they can complete a rewiring job in 3 hours. How long would it take each of them working alone? (Show Your Solution)
Answer by stanbon(57347)

An electrician works thrice as fast as his apprentice. Together they can complete a rewiring job in 3 hours. How long would it take each of them working alone? (Show Your Solution)
----
apprentice rate = 1/3x job/hr
boss's rate = 1/x job/hr
together rate = 1/3 job/hr
----
Equation:
rate + rate = together rate
1/3x + 1/x = 1/3
---
1 + 3 = x
x = 4 hrs (time for the boss)
3x = 12 hrs (time for the apprentice)
---------------------
Cheers,
stan H.